A 50 meter by 120 meter park consists of a rectangular lawn surrounded by a path of uniform width. Find the dimensions of the lawn if its area is the same as the area of the path.
Answer by edjones(8007)

50*120=6000 area of (path + lawn)
area of lawn=3000
(120-x)(50-x)=3000
x^2-170x+6000=3000
x^2-170x+3000=0
(x-20)(x-150)=0
x=20 m twice the width of the path
100 m * 30 m dimensions of the lawn.
.
Ed